I have 600 meters of fabric, 3 products with 3 different yields, 1st product yields 1.5 per meter, 2nd product yields 2 per meter, 3rd yields 3 per meter need to produce 3 equal amounts of product. I want this in an excel spread sheet and cannot solve. This is a real life question no textbook
Answer by scott8148(6628)

you have the total length of fabric, so you neeed the lengths of the products

"1st product yields 1.5 per meter" ___ Lf = 1 / 1.5 = 2 / 3

"2nd product yields 2 per meter" ___ Ls = 1 / 2

"3rd yields 3 per meter" ___ Lt = 1 / 3

let x = "amount of each product"

x(Lf) + x(Ls) + x(Lt) = 600

2/3 x + 1/2 x + 1/3 x = 600

multiplying by 6 (Lowest Common Denominator) ___ 4x + 3x + 2x = 3600 ___ 9x = 3600

dividing by 9 ___ x = 400